A 9.119.11 g sample of calcium sulfide was decomposed into its constituent elements, producing 5.075.07 g of calcium and 4.044.04 g of sulfur.

Which of the statements are consistent with the law of constant composition (definite proportions)?

-The mass ratio of CaCa to SS in every sample of calcium sulfide is 1.25.

-The mass percentage of calcium plus the mass percentage of sulfur in every sample of calcium sulfide equals 100%.

-Every sample of calcium sulfide will have 55.6% mass of calcium.

-Every sample of calcium sulfide will have 5.07 g of calcium.

-The ratio of calcium to sulfur will vary based on how the sample was prepared.
Answer #1

Law of constant composition : The relative amount of each element in a particular compound remains same irrespective of source or how much compound is prepared.

Initial mass of compound calcium sulfide = 9,119.11 g

Mass of calcium product = 5,075.07 g

Mass of sulphur product = 4,044.04 g

a) The mass ratio of CaCa to SS in every sample of calcium sulfide is 1.25 = 5,075.07 g/4,044.04 g = 1.25 is a consistent statement.

Molar mass of Calcium = 40 g

Molar mass of sulphur = 32 g

Mass ratio of Ca/S = 40/32 = 1.25

b) The mass percentage of calcium plus the mass percentage of sulfur in every sample of calcium sulfide equals 100%.

5075.07 g + 4044.04 g = 9119.11 g = mass of initial compound. So, it is consistent

c) Every sample of calcium sulfide will have 55.6% mass of calcium

5075.07 g/9119.11 g ) x 100 = 55.6 %. So, it is consistent

d) Every sample of calcium sulfide will have 5.07 g of calcium. It is not consistent

e) The ratio of calcium to sulfur will vary based on how the sample was prepared is not consistent based on the definition.